# MoLoRec: A Generalizable and Efficient Framework for LLM-Based Recommendation

### v1

@@ Abstract.
Large Language Models (LLMs) have achieved remarkable success in recent years, owing to their impressive generalization capabilities and rich world knowledge. To capitalize on the potential of using LLMs as recommender systems, mainstream approaches typically focus on two paradigms. The first paradigm designs multi-domain or multi-task instruction data for generalizable recommendation, so as to align LLMs with general recommendation areas and deal with cold-start recommendation. The second paradigm enhances domain-specific recommendation tasks with parameter-efficient fine-tuning techniques, in order to improve models under the warm recommendation scenarios. While most previous works treat these two paradigms separately, we argue that they have complementary advantages, and combining them together would be helpful.
To that end, in this paper, we propose a generalizable and efficient LLM-based recommendation framework MoLoRec. Our approach starts by parameter-efficient fine-tuning a domain-general module with general recommendation instruction data, to align LLM with recommendation knowledge. Then, given users’ behavior of a specific domain, we construct a domain-specific instruction dataset and apply efficient fine-tuning to the pre-trained LLM. After that, we provide approaches to integrate the above domain-general part and domain-specific part with parameters mixture. Please note that, MoLoRec is efficient with plug and play, as the domain-general module is trained only once, and any domain-specific plug-in can be efficiently merged with only domain-specific fine-tuning. Extensive experiments on multiple datasets under both warm and cold-start recommendation scenarios validate the effectiveness and generality of the proposed MoLoRec. Codes are available at https://anonymous.4open.science/r/MoLoRec.
1. Introduction
@@
Large Language Models (LLMs) have demonstrated significant success across diverse fields (Zhao et al., 2023), driven by their emergent capabilities (Dong et al., 2024; Huang and Chang, 2023) such as world knowledge, language understanding, and complex reasoning. Recently, LLMs have introduced transformative advancements to recommendation tasks. Notably, LLMs have shown potential in capturing nuanced item semantics (Xi et al., 2024), understanding diverse user interests (Deng et al., 2022), and unifying various recommendation tasks (Geng et al., 2022). These advancements highlight the promise of utilizing LLMs as recommender systems, positioning LLM-based recommendations as a compelling area for further exploration.
@@
In the field of LLM-based recommendation research, the emergence of ChatGPT and its remarkable reasoning capabilities have catalyzed early studies (Dai et al., 2023; Sanner et al., 2023; Wang et al., 2023). These works focus on the zero-shot/few-shot recommendation potential of LLMs through in-context learning (Dong et al., 2024). However, the intrinsic gap between the pre-training general text corpus of LLMs and the requirements of recommendation tasks results in suboptimal performance when relying solely on in-context learning. Consequently, the key to developing an effective LLM-based recommender system lies in bridging this gap, enabling the model to truly ”understand” how to recommend. To address this challenge, researchers have proposed a variety of approaches. We classify them into two paradigms, each tackling the problem from a distinct perspective. As shown in Figure 1(a), the first one is summarized as the breadth-oriented paradigm. These works integrate multi-domain (Tang et al., 2024) or multi-task (Geng et al., 2022; Zhang et al., 2024; Cui et al., 2022) recommendation data to construct extensive recommendation world knowledge, paving the way for developing a generalizable LLM-based recommender. The key focus of this paradigm is the integration of multi-source data to build instruction-tuning datasets (Peng et al., 2024; Jin et al., 2023) and the design of instruction templates (Zhang et al., 2024; Geng et al., 2022) tailored to various tasks. The second paradigm is termed the depth-oriented paradigm, illustrated in Figure 1(b). This line of research seeks to enable LLMs to deeply comprehend recommendation tasks within specific domains. Key areas of focus include: the in-depth extraction of domain-specific recommendation knowledge, such as collaborative filtering information (Lin et al., 2024b; Kim et al., 2024a; Liao et al., 2024; Ren et al., 2024; Kong et al., 2024), and the development of efficient and effective alignment methods between LLMs and recommendation tasks. Specifically, compared with enormous parameters in LLMs, downstream tasks do not have sufficient data for tuning all parameters. Therefore, parameter-efficient fine-tuning methods become optimal for applying LLMs, in which lightweight Low-Rank Adapter (LoRA) is one representative work (Hu et al., 2022). By borrowing ideas of LLMs, these methods include leveraging (Bao et al., 2023) or enhancing (Kong et al., 2024) LoRA fine-tuning techniques and designing data-efficient fine-tuning strategies (Lin et al., 2024c).
@@
These works make significant advancements in recommendation research. Nevertheless, we argue that these two paradigms have complementary advantages. The first paradigm masters general recommendation knowledge and can be well generalized to various recommendation scenarios. The second paradigm learns each user’s unique preference and is suitable for the warm-start recommendation of a specific domain. However, the specific recommendation domain performance could not be easily transferred to other domains. In fact, both generalizable recommendation knowledge and efficient domain-specific understanding are essential for recommender systems. Relying solely on one aspect risks falling short in addressing the diverse challenges in real-world recommendation scenarios. The breadth-oriented paradigm may underperform in specific domains. Conversely, the depth-oriented paradigm struggles with distribution shifts between training and test data. It faces challenges when new users or items appear or when training data are sparse.
@@
To this end, we investigate how to integrate the advantages of both paradigms to simultaneously enhance the model’s generalization ability and domain-specific performance. The task creates significant obstacles: (1) Efficiency. Integrating two paradigms may introduce model complexity, and finding an efficient integrating method without excessive computational overhead is a critical challenge. (2) Generalizability. We need to preserve the model’s generalization ability to a large extent, enabling it to quickly scale to new domains, new items, and other new recommendation scenarios. In this paper, we propose a generalizable and efficient recommendation framework named Mixture-of-LoRA Recommendation Framework (MoLoRec). MoLoRec fuses both general recommendation knowledge and domain-specific knowledge with three key stages. Firstly, to align LLM with any recommendation task, MoLoRec constructs a general recommendation instruction dataset from multiple recommendation domains, and fine-tunes LLM to get a domain-general LoRA module. Secondly, to tailor the framework for specific domains, MoLoRec constructs domain-specific instruction datasets derived from the specific domain, and fine-tunes LLM to get a domain-specific LoRA module. After that, MoLoRec performs a highly efficient and effective linear arithmetic operation to merge these LoRA adapters within the weight space, allowing MoLoRec to maintain strong recommendation performance across both specific domains and out-of-distribution scenarios. To further enhance the merging process of LoRA adapters, we also introduce an adaptive LoRA weight merging method guided by entropy minimization during test time. Importantly, our framework is designed for ease of use, allowing for a plug-and-play integration where the domain-general module is trained once, and domain-specific adaptations are incorporated through minimal fine-tuning. Finally, extensive experiments conducted on various datasets demonstrate the effectiveness and versatility of the framework in both warm recommendation scenarios and challenging cold-start scenarios, highlighting its potential for broad application.
2. Preliminary
In this section, we introduce key concepts underpinning our methodology. First, we cover the task formulation and instruction tuning for LLM-based recommender models. Next, we highlight the use of LoRA for parameter-efficient fine-tuning of LLMs.
2.1. LLM-Based Recommender Models
@@
Task Formulation. We mainly focus on the sequential recommendation task, which holds significant practical importance. Let and represent the sets of users and items, respectively. The historical interaction sequence of a user is denoted as , arranged in chronological order, where and . The goal is to predict this user’s next liked item based on the historical interactions.
@@
Instruction Tuning for LLM-Based Recommendation. For LLM-based recommendation, instruction tuning (Wei et al., 2022) is the key step to bridge the gap between the next-word prediction objective of LLMs and the recommendation task (Bao et al., 2023; Kong et al., 2024; Lin et al., 2024c). Formally, instruction tuning involves fine-tuning LLMs using training data organized as explicit instruction pairs . Here, represents a detailed textual instruction that encapsulates the interaction sequences and the recommendation task, while corresponds to the textual description of the predicted item . The instruction fine-tuning process is guided by minimizing the following autoregressive loss function:
@@
where denotes the -th token of , is the token sequence preceding , and is the LLM’s parameters.
2.2. Low-Rank Adaptation (LoRA)
@@
In traditional fine-tuning as described in Eqn. (1), updating all parameters makes the process highly computationally intensive, particularly for LLMs. To address this issue, parameter-efficient methods are designed to fine-tune LLMs while updating only a small subset of parameters. Low-Rank Adaptation (LoRA) (Hu et al., 2022) is the mainstream approach. LoRA addresses this issue by introducing low-rank matrices that are trained alongside the frozen original model weights. This allows the model to adapt to specific tasks by learning a small number of additional parameters, without requiring modifications to the entire model.
@@
3. Methodology
In this section, we propose MoLoRec, a generalizable, effective, and efficient LLM-based recommendation framework. As shown in Figure 2, MoLoRec operates through three key stages. Firstly, we align the LLM with any recommendation task and train a domain-general LoRA module (Section 3.1). Secondly, to adapt the framework to specific domains, we fine-tunes the LLM to obtain domain-specific LoRA modules (Section 3.2). Subsequently, MoLoRec performs an efficient linear arithmetic operation to merge these LoRA adapters within the weight space, enabling it to maintain strong recommendation performance across both specific domains and out-of-distribution scenarios (Section 3.3). Notably, our framework is designed for ease of use, offering plug-and-play integration, where the domain-general module is trained once, and domain-specific adaptations are incorporated with minimal fine-tuning. MoLoRec is capable of significantly enhancing the recommendation performance and generalization across various domains at a relatively low cost.
3.1. Training Generalizable Base Recommender
Pre-trained LLMs often exhibit suboptimal performance when directly applied to recommendation tasks (Bao et al., 2023; Lin et al., 2024a). This limitation arises because their pre-training on general-purpose datasets fails to capture the specialized knowledge required for understanding user preferences, behaviors, and the contextual nuances essential to building effective recommender systems. To bridge this gap, in this subsection, we align the pre-trained LLM with the recommendation task, as shown in Figure 2(a). Specifically, we first construct large-scale instruction data by mixing users’ behaviors from multiple domains. The pre-trained LLM is then fine-tuned on this dataset using the LoRA technique, equipping it with general knowledge to tackle recommendation tasks.
@@ 3.1. Training Generalizable Base Recommender
Instruction Data Construction. Given recommendation domains (i.e., , ,…, ), denotes , , and as the user set, item set and user interaction sequence set of domain , respectively. To provide general user modeling and recommendation knowledge, we combine data from all domains and design instruction templates to convert them into a text format. Note that the choice of recommendation domains and instruction templates can be arbitrary. As illustrated in Figure 2(a), we transform the recommendation data into instruction data , where x and y denote the instruction input and output, respectively. The instruction input includes the task description (which explains the recommendation task), the user’s historical interactions, and the item candidate set, all in natural language. Here, the items are represented by their titles. The candidate set consists of one ground-truth item and some randomly selected negative samples. The instruction output is designed to rank the user’s next most likely products.
@@
Tuning Domain-General LoRA Module. Given the instruction data , we apply LoRA fine-tuning to adapt the pre-trained LLM for general recommendation tasks. The pre-trained model parameters are kept frozen, while trainable low-rank decomposition matrices are introduced into each layer of the Transformer architecture, enabling efficient and lightweight tuning. Formally,
@@
where is the parameters of the pre-trained LLM, and denotes the set of parameters of LoRA fine-tuning. By undergoing this fine-tuning step, is now enriched with extensive general knowledge relevant to the field of recommendations.
3.2. Training Domain-Specific Plugin
@@
Each recommendation domain exhibits unique user behavior patterns, making the acquisition of domain-specific knowledge essential for delivering accurate recommendations. For instance, in the clothing recommendation domain, user behavior is predominantly driven by style preferences. In contrast, when purchasing electronic products, users tend to prioritize compatibility with their existing products. To address these domain-specific needs, we construct an instruction dataset tailored to the specific new recommendation domain and apply LoRA fine-tuning to the pre-trained LLM, resulting in the LoRA modules enhanced with domain-specific expertise. As shown in Figure 2(b), the template for constructing the instruction dataset and LoRA fine-tuning method are similar to those outlined in Section 3.1. We formally define the process of training the domain-specific LoRA module as:
3.3. Mixture-of-LoRA for Plug-and-Play.
@@
After instruction fine-tuning in stages 1 and 2, we obtained a domain-general model that captures general recommendation knowledge and a domain-specific model that incorporates domain-specific insights. In this subsection, we propose integrating these two parts to improve recommendation accuracy and enhance generalization capabilities simultaneously. Natural questions arise: could this goal be achieved by applying traditional ensemble learning methods that combine the outputs of multiple models? Or could we integrate general and domain-specific knowledge by directly performing second-round fine-tuning on the domain-general LoRA module with the domain-specific dataset? Unfortunately, the answer is no. Since LLMs generate natural language text, ensembling their outputs can introduce semantic inconsistencies or ambiguities, while also increasing inference time and GPU memory usage. Meanwhile, performing a second round of fine-tuning risks catastrophic forgetting (Kirkpatrick et al., 2017), causing the model to collapse.
@@
Mixture-of-LoRA. We propose a simple yet effective method called Mixture-of-LoRA, which linearly merges the model parameters of the domain-general LoRA module and the domain-specific LoRA module . Formally, given the domain-general LoRA module and the domain-specific LoRA module , we define the mixture-of-LoRA operator as:
@@
where the coefficients and represents the importance of merging. We constraint and . They can be considered hyperparameters and selected using the validation data. Please note that the mixed maintains the same total number of parameters as one standard LoRA, making our Mixture-of-LoRA method simple, fast, and effective. There is no extra cost at inference time in terms of memory or compute, since we only do element-wise operations on model weights. In addition, the domain-general LoRA module is reusable. When facing a new domain, it is only necessary to retrain a domain-specific LoRA module.
Mixture-of-LoRA is inspired by recent studies (Ilharco et al., 2023; Wortsman et al., 2022) on the linear connectivity of trained models in a full finetuning setting. These studies suggest that parameters of tuned models can be directly added to improve generalization, provided they are initialized from the same pre-trained model checkpoint. Specifically, as shown in Figure 3, recent research (Ilharco et al., 2023) defines the concept of ”task vector”. A task vector specifies a direction in the weight space of a pre-trained model , such that movement in that direction improves performance on the task. It is built by subtracting the weights of a pre-trained model () from the weights of the same model after fine-tuning () on a task. Adding task vectors together can improve performance on multiple tasks at once. The underlying hypothesis is that two models finetuned from the same pre-trained checkpoint often lie in the same error basin (Neyshabur et al., 2020; Zhang et al., 2023), and thus the parameters could be directly added. Extending this property to the context of LoRA, we hypothesize that LoRA modules can also be linearly combined. This is because a LoRA module can be considered as the difference between a fine-tuned LLM and its pre-trained counterpart, making it analogous to a task vector.
@@
Entropy-Guided Adaptive Mixture-of-LoRA. As shown in Figure 2(c), in this subsection, we provide an efficient and automatic way to better choose mixture coefficients and . As discussed in the previous subsection, and can be chosen by employing the grid-search in the validation data. Nevertheless, (1) it is still lacking a guiding principle. (2) When the distribution of the inference data differs significantly from that of the validation set, the chosen coefficients may perform poorly. To this end, we introduce entropy minimization on the unlabeled test samples an optimization surrogate objective to update and . Specifically, the Shannon entropy (Shannon, 1948) is a well-known measure of uncertainty. For a sample , the predicted output of a neural network is , the Shannon entropy is calculated as , where denotes the probability that the input is predicted to be the -th class. Lower entropy indicates that the model has lower uncertainty about its predictions, meaning the model is more confident in its outputs. Therefore, the intuition behind our method is that the good coefficients and for the test inputs should make the mixed model more confident in its prediction, that is, it should lead to lower model entropy over the input (Wang et al., 2021a, b; Yang et al., 2024b). Formally, we collect a set of unlabeled test samples , i.e., some instruction inputs in the test time. We fix the , , , and using the following entropy minimization loss to update coefficients and :
@@
For the LLM, the output of is a sentence. Since our instruction is to select a title from a given candidate set, the first few tokens output by the model are more important because after deciding on them, the subsequent tokens are more certain. So in practice, we only calculate the average entropy of the first three tokens in the sentence to represent . Besides, we do not need all test data to be available. Even if only 50 unlabeled tests are available, our method can have significant performance improvements.
3.4. Discussion

### v2

# RecCocktail: A Generalizable and Efficient Framework for LLM-Based Recommendation

@@ Abstract
Large Language Models (LLMs) have achieved remarkable success in recent years, owing to their impressive generalization capabilities and rich world knowledge. To capitalize on the potential of using LLMs as recommender systems, mainstream approaches typically focus on two paradigms. The first paradigm designs multi-domain or multi-task instruction data for generalizable recommendation, so as to align LLMs with general recommendation areas and deal with cold-start recommendation. The second paradigm focuses on enhancing domain-specific recommendation tasks, improving performance in warm recommendation scenarios. While most previous works treat these two paradigms separately, we argue that they have complementary advantages, and combining them can yield better results. In this paper, we propose a generalizable and efficient LLM-based recommendation framework RecCocktail. Our approach begins with fine-tuning a “base spirit” LoRA module using domain-general recommendation instruction data to align LLM with recommendation knowledge. Next, given users’ behavior of a specific domain, we construct a domain-specific “ingredient” LoRA module. We then provide an entropy-guided adaptive merging method to mix the “base spirit” and the “ingredient” in the weight space. Please note that, RecCocktail combines the advantages of the existing two paradigms without introducing additional time or space overhead during the inference phase. Moreover, RecCocktail is efficient with plug and play, as the “base spirit” LoRA is trained only once, and any domain-specific “ingredient” can be efficiently mixed with only domain-specific fine-tuning. Extensive experiments on multiple datasets under both warm and cold-start recommendation scenarios validate the effectiveness and generality of the proposed RecCocktail.
Code — https://anonymous.4open.science/r/RecCocktail
1 Introduction
@@
Large Language Models (LLMs) have demonstrated significant success across diverse fields (Zhao et al. 2023), driven by their emergent capabilities (Dong et al. 2024; Huang and Chang 2023) such as world knowledge, language understanding, and complex reasoning. Recently, LLMs have introduced transformative advancements to recommendation tasks (Xi et al. 2024; Deng et al. 2022; Geng et al. 2022).
Along this line, the emergence of ChatGPT and its remarkable reasoning capabilities have catalyzed early studies (Dai et al. 2023; Sanner et al. 2023; Wang et al. 2023). These works focus on the zero-shot/few-shot recommendation potential of LLMs through in-context learning (Dong et al. 2024). However, the intrinsic gap between the pre-training general text corpus of LLMs and the requirements of recommendation tasks results in suboptimal performance when relying solely on in-context learning. Consequently, the key to developing an effective LLM-based recommender system lies in bridging this gap, enabling the model to truly “understand” how to recommend.
@@
To address this challenge, researchers have proposed a variety of approaches. We classify them into two paradigms, each tackling the problem from a distinct perspective. As shown in Figure 1(1), the first one is summarized as the breadth-oriented paradigm. These works integrate multi-domain (Tang et al. 2024) or multi-task (Geng et al. 2022; Zhang et al. 2024; Cui et al. 2022) recommendation data to construct extensive recommendation world knowledge, paving the way for developing a generalizable LLM-based recommender. The key focus of this paradigm is the integration of multi-source data to build instruction-tuning datasets (Peng et al. 2024; Jin et al. 2023) and the design of instruction templates (Zhang et al. 2024; Geng et al. 2022) tailored to various tasks. The second paradigm is termed the depth-oriented paradigm, illustrated in Figure 1(2). This line of research seeks to enable LLMs to deeply comprehend recommendation tasks within specific domains. Key areas of focus include: the in-depth extraction of domain-specific recommendation knowledge, such as collaborative filtering information (Lin et al. 2024a; Kim et al. 2024a; Liao et al. 2024; Ren et al. 2024; Kong et al. 2024), and the development of efficient and effective alignment methods between LLMs and recommendation tasks. Specifically, compared with enormous parameters in LLMs, downstream tasks do not have sufficient data for tuning all parameters. Therefore, parameter-efficient fine-tuning methods become optimal for applying LLMs, in which lightweight Low-Rank Adapter (LoRA) is one representative work (Hu et al. 2022). By borrowing ideas of LLMs, these methods include leveraging (Bao et al. 2023) or enhancing (Kong et al. 2024) LoRA fine-tuning techniques and designing data-efficient fine-tuning strategies (Lin et al. 2024b).
@@
These works make significant advancements in recommendation research. Nevertheless, we argue that these two paradigms have complementary advantages. Both generalizable recommendation knowledge and efficient domain-specific understanding are essential for recommender systems. Relying solely on one aspect risks falling short in addressing the diverse challenges in real-world recommendation scenarios. The breadth-oriented paradigm may underperform in specific domains. Conversely, the depth-oriented paradigm struggles with distribution shifts between training and test data. It faces challenges when new users or items appear or when training data are sparse.
@@
To this end, we investigate how to integrate the advantages of both paradigms to simultaneously enhance the model’s generalization ability and domain-specific performance. The task creates significant obstacles: (1) Efficiency. Integrating two paradigms may introduce model complexity, and finding an efficient integrating method without excessive computational overhead is a critical challenge. (2) Generalizability. We need to preserve the model’s generalization ability to a large extent, enabling it to quickly scale to new domains, new items, and other new recommendation scenarios. In this paper, we propose a generalizable and efficient recommendation framework named RecCocktail, inspired by the cocktail preparation process, as shown in Figure 1(3). Specifically, to align LLM with any recommendation task, RecCocktail constructs a general recommendation instruction dataset from multiple recommendation domains, and fine-tunes LLM to get a domain-general LoRA module as “base spirit”. Secondly, to tailor the framework for specific domains, RecCocktail constructs domain-specific instruction datasets, and fine-tunes LLM to get a domain-specific LoRA module as “ingredient”. After that, RecCocktail performs a highly efficient and effective linear arithmetic operation to merge the “base spirit” and the “ingredient” LoRA within the weight space, allowing RecCocktail to maintain strong recommendation performance across both specific domains and out-of-distribution scenarios. To further enhance the merging process, we also introduce an adaptive merging method guided by entropy minimization during test time. Importantly, RecCocktail does not introduce additional time or space overhead during the inference phase. Furthermore, RecCocktail is designed for ease of use, allowing for a plug-and-play integration where the “base spirit” is trained once, and domain-specific “ingredients” are incorporated through minimal fine-tuning. Finally, extensive experiments conducted on various datasets demonstrate the effectiveness and generalizability of the framework in multiple recommendation scenarios, highlighting its potential for broad application.
2 Preliminary
2.1 LLM-Based Recommendation
@@
Task Formulation. Let and represent the sets of users and items, respectively. The historical interaction sequence of a user is denoted as , arranged in chronological order, where and . The goal is to predict this user’s next liked item based on the historical interactions.
@@
Instruction Tuning for LLM-Based Recommendation. For LLM-based recommendation, instruction tuning (Wei et al. 2022) is the key step to bridge the gap between the next-word prediction objective of LLMs and the recommendation task (Bao et al. 2023; Kong et al. 2024; Lin et al. 2024b). Formally, instruction tuning involves fine-tuning LLMs using training data organized as explicit instruction pairs . Here, represents a detailed textual instruction that encapsulates the interaction sequences and the recommendation task, while corresponds to the textual description of the predicted item . The instruction fine-tuning process is guided by minimizing the following autoregressive loss function:
@@
where denotes the -th token of the output sequence , is the token sequence preceding , and is the LLM’s model parameters.
2.2 Instruction Tuning with LoRA
@@
In traditional fine-tuning as described in Eqn. (1), updating all parameters makes the process highly computationally intensive, particularly for LLMs. To address this issue, parameter-efficient methods are designed to fine-tune LLMs while updating only a small subset of parameters. Low-Rank Adaptation (LoRA) (Hu et al. 2022) is the mainstream approach. LoRA addresses this issue by introducing low-rank matrices that are trained alongside the frozen original model weights. This allows the model to adapt to specific tasks by learning a small number of additional parameters, without requiring modifications to the entire model.
@@
3 Methodology
In this section, we propose RecCocktail, a generalizable, effective, and efficient LLM-based recommendation framework. As shown in Figure 2, RecCocktail operates through three key stages. We start by preparing the base spirit. We align the LLM with any recommendation task and fine-tune a domain-general LoRA module as the base spirit (Section 3.1). Secondly, to adapt the framework to a specific domain, we fine-tune the domain-specific LoRA module (Section 3.2) as an ingredient. Subsequently, RecCocktail performs an efficient linear arithmetic operation to merge the base spirit LoRA and the ingredient LoRA into a cocktail within the weight space (Section 3.3). We further provide a test-time entropy-guided adaptive merging method to quickly construct cocktail LoRA tailored to different inference scenarios. The merged cocktail LoRA collaborative enhances performance across both specific domains and out-of-distribution scenarios without introducing additional time or space overhead.
3.1 Constructing Generalizable Base Spirit
In this subsection, we construct a generalizable base spirit equipped with general recommendation knowledge. As illustrated in Figure 2(a), we first create a large-scale instruction dataset by aggregating user behavior data from multiple domains. We then fine-tune a pre-trained LLM on this dataset using the LoRA technique, resulting in a domain-general LoRA module, which serves as our base spirit.
@@ 3.1 Constructing Generalizable Base Spirit
Instruction Dataset Construction. Given recommendation domains (i.e., , ,…, ), let , , and denote the user set, item set, and user interaction sequence set of domain , respectively. To provide general user modeling and recommendation knowledge, we aggregate data from all domains and design instruction templates to convert them into text format. Note that both the choice of recommendation domains and the design of instruction templates are flexible. In this paper, we adopt the template shown in Figure 2(a). We transform the multi-domain recommendation data into an instruction dataset , where x and y represent the instruction input and output, respectively. The input a task description, the user’s historical interactions, and a set of candidate items, all expressed in text. Each item is represented by its title. The candidate set includes one ground-truth item and several randomly sampled negative samples. The output is designed to rank the user’s next most likely item.
@@
Tuning Domain-General LoRA Module. Given the instruction dataset , we apply LoRA fine-tuning to adapt the pre-trained LLM for general recommendation tasks. The pre-trained model parameters are kept frozen, while trainable low-rank decomposition matrices are introduced into each layer of the Transformer architecture, enabling efficient and lightweight tuning. Formally,
@@
where is the parameters of the pre-trained LLM, and denotes the set of parameters of LoRA fine-tuning. By undergoing this fine-tuning step, is now enriched with extensive general recommendation knowledge.
3.2 Constructing Domain-Specific Ingredient
@@
Each recommendation domain exhibits unique user behavior patterns, making the acquisition of domain-specific knowledge essential for delivering accurate recommendations. To address these domain-specific characteristics, we construct an instruction dataset tailored to the target domain , and fine-tune the pre-trained LLM using the LoRA technique. The domain-specific LoRA module serves as a ingredient of a cocktail. As shown in Figure 2(b), the instruction template and LoRA fine-tuning procedure follow a similar approach to that described in Section 3.1. Formally, the training objective for the domain-specific LoRA is defined as:
3.3 LoRA Cocktail for Plug-and-Play
@@
After preparing the base spirit and the ingredient, we propose integrating these two parts to improve recommendation accuracy and enhance generalization capabilities simultaneously. Natural questions arise: could this goal be achieved by applying traditional ensemble learning methods that combine the outputs of multiple models? Or could we integrate general and domain-specific knowledge by directly performing second-round fine-tuning on the domain-general LoRA module with the domain-specific dataset? Unfortunately, the answer is no. Since LLMs generate natural language text, ensembling their outputs can introduce semantic inconsistencies or ambiguities, while also increasing inference time and GPU memory usage. Meanwhile, performing a second round of fine-tuning risks catastrophic forgetting (Kirkpatrick et al. 2017), causing the model to collapse.
@@
LoRA Cocktail. We propose a simple yet effective method called LoRA cocktail, which linearly merges the model parameters of the domain-general base spirit LoRA and the domain-specific ingredient LoRA . We draw an analogy to cocktail making, where mixing a base spirit with an additional ingredient results in a drink that combines the flavors of both parts. Formally, given the base spirit and a ingredient , we define the LoRA cocktail operator as:
@@
where the coefficients and represents the importance of merging. We constraint and . They can be considered hyperparameters and selected using the validation data. Please note that the mixed maintains the same total number of parameters as one standard LoRA, making our LoRA Cocktail method simple, fast, and effective. There is no extra cost at inference time in terms of memory or compute, since we only do element-wise operations on model weights. In addition, the domain-general LoRA module is reusable. When facing a new domain, it is only necessary to retrain a domain-specific LoRA.
LoRA cocktail is the weight merging of the domain-general LoRA module and the domain-specific one. Here we explain how it works as shown in Figure 3. We find that after fine-tuning, a LoRA module can be interpreted as a task-specific weight update vector: , which defines a direction in the pre-trained model’s weight space. Shifting the model’s weights along this direction enhances its performance on the corresponding task. Intuitively, a LoRA vector encodes all of the information needed to solve a task that is learned via fine-tuning. Adding two directions allows the model to move in a way that incorporates both general knowledge and domain-specific knowledge. This idea aligns with the concept of “task vectors” proposed in recent work (Ilharco et al. 2023), where weight differences from pre-trained to fine-tuned models are shown to encode task-specific information. These vectors, residing in the same parameter space, can be combined to equip the model with multiple capabilities. This is further supported by findings that models fine-tuned from the same initialization often lie in the same error basin (Neyshabur, Sedghi, and Zhang 2020; Zhang et al. 2023), making such linear merging feasible and effective. In this light, the LoRA cocktail approach leverages LoRA as task vectors, enabling efficient knowledge integration through simple vector addition.
@@
Entropy-Guided Adaptive Merging. As shown in Figure 2(c), we provide an efficient and automatic way to better choose merging coefficients and . As discussed in the previous subsection, and can be chosen by employing the grid-search in the validation data. Nevertheless, (1) it is still lacking a guiding principle. (2) When the distribution of the inference data differs significantly from that of the validation set, the chosen coefficients may perform poorly. To this end, we introduce entropy minimization on the unlabeled test samples as an optimization surrogate objective to update and . Specifically, the Shannon Entropy (Shannon 1948) is a well-known measure of uncertainty. For a sample , the predicted output of a neural network is , the Shannon entropy is calculated as , where denotes the probability that the input is predicted to be the -th class. Lower entropy indicates that the model has lower uncertainty about its predictions, meaning the model is more confident in its outputs. Therefore, the intuition behind our method is that the good coefficients and for the test inputs should make the mixed model more confident in its prediction, that is, it should lead to lower model entropy over the input (Wang et al. 2021a, b; Yang et al. 2024c). Formally, we collect a set of unlabeled test samples , i.e., some instruction inputs in the test time. We fix the , , , and using the following entropy minimization loss to update coefficients and :
@@
For the LLM, the output of is a sentence. Since our instruction is to select a title from a given candidate set, the first few tokens output by the model are more important because after deciding on them, the subsequent tokens are more certain. So in practice, we can only calculate the average entropy of the first few tokens in the sentence to represent . Good performance can be achieved with just 3 tokens. Besides, we do not need all test data to be available. Even with only 50 unlabeled tests data, our method can have significant performance improvements.
3.4 Discussion